OpenStudy (anonymous):

In a simple physical apparatus, the distance x cm and y cm are related by the equation 1/x + 1/y = 1/5 . if y is increasing at a rate of 16 cm per minute, calculate the rate at which x is changing when x=15cm. [Note: apparatus: Equipment designed to serve a specific function] [Oct08: -64cm/min]

OpenStudy (anonymous):

how to solve it?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

hmm.. i got dx/dt = -4cm/min but i dont know if its correct or not

OpenStudy (anonymous):

if its correct then tell me and ill post up the worked solution

OpenStudy (anonymous):

it's ok coz smetimes the ans given is incorrect.

OpenStudy (anonymous):

dx/dt = -25(x-5)^2 times 16, then sub x=15 and u get -4cm/sec

OpenStudy (dumbcow):

no its -64...hold on

OpenStudy (anonymous):

ok dumbcow

OpenStudy (anonymous):

ok i got it, dx/dt = (16*(x-5)^2)/-25 sub x=15 and you get -64cm/min

myininaya (myininaya):

i got x'=-576 unless i made a mistake

OpenStudy (anonymous):

thegreatsaiyaman: can u explaim to me how u get the ans?

OpenStudy (dumbcow):

\[\frac{dx}{dt} = \frac{dy}{dt}*\frac{dx}{dy}\] if you solve the equation for x \[x = \frac{5y}{y-5}\] \[\frac{dx}{dy} = -\frac{25}{(y-5)^{2}}\] \[\frac{dx}{dt} = -\frac{16*25}{(y-5)^{2}}\] Now they don't tell us what y is, only that x=15 so substitute 15 in to find the corresponding y-value \[\frac{1}{y} = \frac{1}{5}-\frac{1}{15} = \frac{2}{15} \rightarrow y = \frac{15}{2}\] \[\frac{dx}{dt} = -\frac{16*25}{(7.5-5)^{2}} = -64\]

OpenStudy (anonymous):

*can u explain

OpenStudy (anonymous):

dumbcows got it, i just dont know how to use these symbols on this website =.=

myininaya (myininaya):

omg darn it i made a mistake

myininaya (myininaya):

i forgot to write the 1 in front of 5 lol

OpenStudy (anonymous):

huhu.. thanks all~

myininaya (myininaya):

15/2 not 5/2 is y

myininaya (myininaya):

yes i get x'=-64 as well lol

OpenStudy (anonymous):

<3 calculas

myininaya (myininaya):

hey bell if you want to see the way i did it here you go

OpenStudy (anonymous):

how 5y become 25?

OpenStudy (dumbcow):

from taking the derivative. using the quotient rule \[\frac{f'g - fg'}{g^{2}}\]

OpenStudy (anonymous):

ouh... ok..

OpenStudy (anonymous):

thanks again to all~